OpenStudy (anonymous):

The length of a rectangle is 3m more than twice its width, and the area of the rectangle is 77m^2 . Find the dimensions of the rectangle.

OpenStudy (anonymous):

that would be the formula to find the length and width?

hero (hero):

Well, it gave the area, so we have to use the formula for area of a rectangle to find it A = xy

hero (hero):

2x + 3 = length x = width A = x(2x+3) 77 = 2x^2 + 3x 2x^2+3x - 77 = 0 Now it's correct :D

OpenStudy (anonymous):

so to get the length and width have to solve the equation right?

hero (hero):

Yes 2x^2+3x - 77 = 0 2x^2 + 14x - 11x - 77 = 0 2x(x+7)-11(x+7) = 0 (x+7)(2x-11) = 0 x = -7 x = 11/2 or 5.5 2x + 3 = 14

hero (hero):

5.5 x 14 = 77

hero (hero):

By the way x = -7 is not a solution

OpenStudy (anonymous):

so then if the width is -7 ..?

OpenStudy (anonymous):

and the length is 14

hero (hero):

No. If you notice, there are two solutions for x. I already told you that x = -7 is not a solution.

OpenStudy (anonymous):

then the length is 14 "?